What is the pH of a 0.84 M solution of NaCN (Ka of HCN = 6.2x10^-10)
Answer #1

CN^{-}+H_{2}O\rightleftharpoons HCN+OH^{-}

0.84 0 0

0.84-x x x

K_{a}\cdot K_{b}=K_{w}

K_{b}=\frac{10^{-14}}{6.2\times10^{-10}}

K_{b}=1.613\times10^{-5}

\frac{[HCN][OH^{-}]}{[CN^{-}]}=1.613\times10^{-5}

\frac{x^{2}}{0.84-x}=1.613\times10^{-5}

x=0.0037

[OH^{-}]=0.0037\, M

[H^{+}]=\frac{10^{-14}}{0.0037}

[H^{+}]=2.7\times10^{-12}M

pH=-log[H^{+}]=-log(2.7\times10^{-12})

pH=11.57